Hint: The area of a circle is calculated as πr2. A square sheet of metal has sides 2p. The circle fits perfectly into the square:


1
Expert's answer
2022-03-01T10:39:00-0500

The radius of the circle is "2p\/2=p." Then the area of a circle is calculated as "\pi p^2."

The area of the square is "(2p)^2=4p^2."

The area between the square and the circle is


"4p^2-\pi p^2=(4-\pi)p^2."